1. A metallic bar 250mm long is rectangular, 30mm x 75mm in section. If is subjected to a uniform stress of 210N/mm^2 along its length, find the changes in dimensions of the bar. E = 205KN/mm^2 and poison's ratio is 0.3


2. A rapper rod having end diameters 8cm and 2.5cm Long. Determine it's extension under a tensile load of 1200kg. E = 2*10^6kg/cm^2


1
Expert's answer
2021-02-02T01:15:27-0500

1.



2.

given information

load(p)=1200Kg

diameter =8cm

E=2*106kg/cm2

C.Area=50.26

L=2.5cm

elongation=PL/AE

1200*2.5/2000000*50.26

=2.984*105 cm